OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (anonymous):

eq. of line parallel to 2x+3y=5 is 2x+3y=c Because it passes through (9,-3) 2*9+3*-3=c 18-9=c, c=9 equation of reqd. line is 2x+3y=9